OpenStudy (anonymous):

If F(x) = 9x, which of the following is the inverse of F(x)? A. F-1(x) = x – 9 B. F-1(x) = C. F-1(x) = D. F-1(x) = x + 9

OpenStudy (anonymous):

what is B. and C.

OpenStudy (anonymous):

A. F-1(x) = x – 9 B. F-1(x) = x/9 C. F-1(x) = 9/x D. F-1(x) = x + 9

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

it would be B.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome